# Estimating Body and Hand Motion in an Ego-sensed World

### v2

@@ 4.1.1 Invariant conditioning evaluation
Figure 5: Egocentric human motion estimation for a squatting sequence. The contents of this figure mirror Figure 4, but use spatial shifts to visualize different timesteps within a sequence.
4.1.2 Comparisons against baselines
To further study EgoAllo’s body estimation quality, we compare against three baselines. (1) NoShape. First, NoShape refers to a variation of EgoAllo that turns off shape estimation, and thus cannot estimate the wearer’s height. (2) EgoEgo. We also compare against the human motion diffusion model from EgoEgo [46]. This is similar to EgoAllo, but considers only the SMPL “mean” body shape and uses canonicalized coordinates for conditioning and as model output. (3) VAE+Opt. Finally, we compare against an approach based on the SLAHMR [99] framework for human motion estimation from exocentric video. A key advantage of SLAHMR is that it uses an unconditional motion prior [71] in an optimization framework. It can therefore be adapted to new settings without re-training—we keep the same body pose and shape variables as the original pipeline, but replace the exocentric keypoint [68] cost with an egocentric CPF pose alignment cost.
EgoAllo improves body motion estimates. We report metrics in Table 2 and visualize example outputs in Figures 4 and 5. We find that EgoAllo enables significant estimation improvements across all datasets, including 20$\sim$30accuracy improvements over EgoEgo for both shorter and longer evaluation sequences. We found shape estimation critical for producing metric-scale, grounded estimates of human body motion, with the head aligned to input SLAM poses and the feet planted on the observed ground plane. This is evident in qualitative results, improved grounding metrics, and in the 6$\sim$7MPJPE gap between EgoAllo and the NoShape ablation.
VAE optimization converges poorly. Optimization-based estimation approaches have been effective for settings with keypoint costs [71, 99], but we found convergence difficult in our less constrained setting. In Table 2, we observe poor generalization: VAE+Opt performs competitively on the AMASS test set, but performance deteriorates dramatically when evaluating on RICH or Aria Digital Twins. VAE+Opt outputs in Figure 4 also look overly smoothed, without the same expressiveness as the conditional predictions of EgoAllo or EgoEgo [46]. This highlights the advantage of using a conditional diffusion model problem for this estimation problem.
4.2 Hand estimation
To evaluate evaluate hands estimated by EgoAllo, we run HaMeR on the segment of the EgoExo4D [16] validation set that is labeled with 3D hand pose keypoints. We quantitatively compare four hand estimation methods in Table 3. In (1) HaMeR [63], we use HaMeR out-of-the-box on undistorted egocentric RGB images. We do not assume bounding boxes as input; instead, we follow the HaMeR demo code and compute crops using ViTPose [96]. When multiple hands are detected for a single side, we naively take the first one. (2) EgoAllo-Mono refers our full system, which uses the same monocular HaMeR hand estimates for guidance. In (3) EgoAllo-Wrist3D, we augment our system with wrist pose estimates from Project Aria’s Machine Perception Services [82]—unlike HaMeR, which assumes monocular input, this uses a pair of SLAM cameras that are unique to Project Aria. Finally, (4) EgoAllo-NoReproj removes the reprojection term (Equation 14) from EgoAllo-Mono. Instead, hand guidance is done directly using the 3D wrist poses predicted by HaMeR. For fairness across settings, we compute metrics only on timesteps where HaMeR estimates are available.
Quantitative results are provided in Table 3. While HaMeR’s local poses (PA-MPJPE) are slightly better, EgoAllo’s hand-body estimation significantly improves how well hands are estimated in the world coordinate system. Compared to HaMeR, EgoAllo drops MPJPE from $237.90\text{mm}\to 131.45\text{mm}$. Incorporating more accurate wrist pose estimates (EgoAllo-Wrist3D) offers a practical solution for further improvements: $131.45\text{mm}\to 60.08\text{mm}$. Reprojection-based guidance is also more robust: EgoAllo-NoReproj outputs are the worst in both MPJPE and PA-MPJPE.
Qualitatively, we observed that high hand estimation errors in naive monocular estimation with HaMeR are explained by a combination of detection failures and monocular ambiguities. Even when detections succeed, the scale and distance of monocular HaMeR estimates are often incorrect or flicker in between frames. Incorporating these hands via guidance with our diffusion motion prior encourages final outputs that obey the kinematic and smoothness constraints imposed by plausible body motion—we provide examples of HaMeR estimates rendered jointly with EgoAllo outputs in Figure 6, where we note realistic elbow and shoulder poses.
Method MPJPE $\downarrow$ PA-MPJPE $\downarrow$
HaMeR 237.90$\pm$1.89 13.04$\pm$1.89
EgoAllo-Mono 131.45$\pm$0.39 14.71$\pm$0.39
EgoAllo-Wrist3D 60.08$\pm$0.26 14.38$\pm$0.26
EgoAllo-NoReproj 143.20$\pm$0.42 14.75$\pm$0.42
5 Dicussion
Limitations and future work. While the core contributions of EgoAllo are general, the current implementation of our system has a few limitations that we hope to explore in future work. First, diffusion model guidance is a test-time optimization process that depends on hyperparameters and incurs a runtime cost. In the future, it may be possible to bootstrap using outputs from our model to train a feedforward model that avoids this step. Success for hand guidance also still depends on reasonable monocular hand estimates. Estimation can therefore fail as a result of errors like left/right flipping or spurious detections, which we found causes high errors in our Table 3 hand estimation metrics. Finally, we also train only on AMASS [54], which includes floor planes but no detailed scene geometry. As a result, our method will fail if we cannot detect an approximate floor plane; this is most likely in settings like hills or staircases. In the future, we hope to extend our insights to data with more detailed scene information, which concurrent work has highlighted the usefulness of in informing human body estimation [20].
Conclusion. We presented EgoAllo, a system for estimating human motion using sensors from head-mounted devices. Our method takes advantage of motion capture data [54] and an off-the-shelf visual hand estimator [63] to jointly estimate human body pose, height, and hand parameters. EgoAllo highlights the importance of spatial and temporal invariance in conditioning for this problem, while demonstrating how estimated bodies can be used to improve hand estimation.
Acknowledgments. This project was funded in part by NSF:CNS-2235013 and IARPA DOI/IBC No. 140D0423C0035. YM acknowledges support from the joint Simons Foundation-NSF DMS grant #2031899, the ONR grant N00014-22-1-2102, the NSF grant #2402951, and partial support from TBSI, InnoHK, and the University of Hong Kong. JM was supported by ONR MURI N00014-21-1-2801. BY is supported by the National Science Foundation Graduate Research Fellowship Program under Grant DGE 2146752. The authors would also like to thank Hongsuk Choi, Michael Taylor, Tyler Bonnen, Songwei Ge, Chung Min Kim, and Justin Kerr for insightful technical discussion and suggestions, as well as Jiaman Li for helpful answers to questions about EgoEgo.
Appendix A Appendix
A.1 Floor height estimation
@@
A key requirement of our algorithm is SLAM poses that can be situated relative to the floor. While floor heights are provided in our training data, they are not directly available on real-world data. We found that a simple RANSAC-based algorithm works well on real-world data from Project Aria [58]. We filter SLAM points by confidence, then use RANSAC to find a z-value with that best fits a plane. Example floor plane outputs using scenes from the EgoExo4D [16] dataset are shown in Figure 7.
A.2 Sequence length evaluation
At test time, EgoAllo follows MultiDiffusion [3] for extrapolating to sequences that are longer than the 128 that we use for training. To evaluate that this works, we filter out test sequences shorter than 256 frames and then evaluate both EgoAllo and EgoEgo [46] with subsequences of size 32, 128, and 256. We report metrics on these sequences in Table 4. Both EgoAllo and EgoEgo include windowing strategies for handling longer sequences; unlike prior work, however, we find that the EgoAllo system improves even after test set sequence lengths surpass the training set sequence length.
Seqlen 32 128 256
EgoAllo 149.3 130.3 127.9
EgoEgo 187.7 173.8 184.3
Table 4: How does MPJPE change with sequence length?
A.2.1 Biomech57 evaluation details
@@
The majority of our evaluation data (AMASS [54] and RICH [25]) is provided directly using SMPL conventions. Because our model outputs SMPL-H parameters, this makes computation of joint error metrics straightforward.
@@
A.3 Implementation details
We use a transformer [91] architecture with rotary positional embeddings [85] for EgoAllo’s denoising model $\mu_{\theta}(\vec{x}_{n},\vec{c},n)$. Latent representations $\vec{z}_{c}$ for conditioning sequences $\vec{c}$ are computed using six transformer blocks, each containing a self-attention layer followed by a 2-layer MLP. Skip connections are used throughout. The denoised output is produced by using six additional transformer blocks that take $\vec{x}_{n}$ as input, while conditioning on $\vec{z}_{c}$ via cross-attention. All hidden dimensions are set to 512. For guidance, we use infrastructure for sparse nonlinear least squares implemented in [100]. For additional details, we refer to our code release.
A.4 Runtimes
Without guidance, each denoising step takes around 0.05 seconds on an RTX 4090; our experiments use 30 DDIM steps for sampling. The guidance optimizer can typically converge in around 0.2 seconds on GPU. Runtimes are reported on length-128 sequences.
Figure 7: Floor height examples. Point cloud-derived floor height examples on the EgoExo4D dataset.

### v3

@@ A.2.2 Additional qualitative results
Figure A.5: Head pose-conditioned motion prior results for a squatting sequence. Spatial shifts are used to visualize different timesteps within the sequence. Hand observations are not used.
A.3 Implementation Details
A.3.1 Network architecture
EgoAllo uses a transformer [94] architecture with rotary positional embeddings [88] for its denoising model $\mu_{\theta}(\vec{x}_{n},\vec{c},n)$. Sampling is performed by denoising all timesteps within a temporal window in parallel: we do not sample autoregressively and therefore do not use causal masking. Encoder details: latent encodings $\vec{z}_{c}$ are computed as output from conditioning sequences $\vec{c}$ as input using six transformer blocks, each containing a self-attention layer followed by a 2-layer MLP. Decoder details: the denoised output is computed using six additional transformer blocks that take $\vec{x}_{n}$ as input, while conditioning on $\vec{z}_{c}$ via cross-attention. All hidden dimensions are set to 512.
Runtime. For a length-128 sequence, each forward pass through EgoAllo’s denoising network takes 0.05 seconds on a single RTX 4090. Because we use DDIM [86] for sampling, the number of denoising steps for each sample can be chosen to make tradeoffs between sample quality and speed. All experiments in our paper use 30 DDIM steps.
A.3.2 Guidance optimizer
For guidance, we use a Levenberg-Marquardt optimizer implemented in JAX [5]. Levenberg-Marquardt is an iterative nonlinear least squares algorithm, which requires solving a linearized subproblem at each timestep. We compute the Jacobians needed for this as block-sparse matrices for efficiency, and solve the resulting linear subproblems using a Conjugate Gradient optimizer.
Runtime. The guidance optimizer converges in 0.15$\sim$0.2 seconds on an RTX 4090. We compare our LM optimizer against off-the-shelf PyTorch optimizers in Figure A.6.
(a) Costs over time. LM converges significantly faster than off-the-shelf PyTorch optimizers for guidance optimization.
Optimizer Final Cost
Levenberg-Marquardt (ours) 209.05
torch.optim.LBFGS 215.96
torch.optim.Adam (lr=1e-2) 248.01
torch.optim.Adam (lr=1e-1) 1733.30
(b) Final costs. We report the final cost for each method in the plot above.
Figure A.6: Comparing guidance optimizers.
A.3.3 Floor height estimation
@@
One requirements of EgoAllo is SLAM poses that can be situated relative to the floor. While floor heights are provided in our training data, they are not directly available on real-world data. We found that a RANSAC-based algorithm works well on real-world data from Project Aria [61]. We filter SLAM points by confidence, then use RANSAC to find a z-value with that best fits a plane. Example floor plane outputs using scenes from the EgoExo4D [17] dataset are shown in Figure A.7.
Figure A.7: Floor height examples. Point cloud-derived floor height examples on the EgoExo4D dataset.
A.3.4 Biomech57 evaluation details
@@
The majority of our evaluation data (AMASS [56] and RICH [26]) is provided directly using SMPL conventions. Because EgoAllo outputs SMPL-H parameters, this makes computation of joint error metrics straightforward.